Find all solutions to the equation.
2 cosx + 1 = 0
If you could please show all steps, I would appreciate it. Thank you!

Answer by Alan3354(69443)

Find all solutions to the equation.
If you mean
2cos(x) + 1 = 0
Subtract 1, divide by 2.
cos(x) = -1/2
x = 120 ± 2n*360 degs, n = 0,1,2,3...
x = 240 ± 2n*360 degs, n = 0,1,2,3...